### v1

# L3Cube-IndicNews: News-based Short Text and Long Document Classification Datasets in Indic Languages

Abstract
In this work, we introduce L3Cube-IndicNews, a multilingual text classification corpus aimed at curating a high-quality dataset for Indian regional languages, with a specific focus on news headlines and articles. We have centered our work on 10 prominent Indic languages, including Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, and Punjabi. Each of these news datasets comprises 10 or more classes of news articles. L3Cube-IndicNews offers 3 distinct datasets tailored to handle different document lengths that are classified as: Short Headlines Classification (SHC) dataset containing the news headline and news category, Long Document Classification (LDC) dataset containing the whole news article and the news category, and Long Paragraph Classification (LPC) containing sub-articles of the news and the news category. We maintain consistent labeling across all 3 datasets for in-depth length-based analysis. We evaluate each of these Indic language datasets using 4 different models including monolingual BERT, multilingual Indic Sentence BERT (IndicSBERT), and IndicBERT. This research contributes significantly to expanding the pool of available text classification datasets and also makes it possible to develop topic classification models for Indian regional languages. This also serves as an excellent resource for cross-lingual analysis owing to the high overlap of labels among languages. The datasets and models are shared publicly at https://github.com/l3cube-pune/indic-nlp.
Keywords: Low Resource Languages, Indic Languages, Web Scraping, News Article Datasets, BERT, Short Text Classification, Long Documents.
L3Cube-IndicNews: News-based Short Text and Long Document Classification Datasets in Indic Languages
[…]
1 Introduction
India boasts a rich linguistic diversity, with over 700 languages spoken, out of which 22 are officially recognized. Some of the primarily spoken languages include Hindi, Bengali, Marathi, Telugu, Tamil, Gujarati, Kannada, Odia, Malayalam, and Punjabi. Despite their widespread use, there’s a notable scarcity of comprehensive Indic language datasets, primarily due to their low-resource status and linguistic complexity Patil and Patil (2017).
These Indic languages are widely spoken all over India and have abundant data available on news websites, and social media but it’s difficult to find a labeled dataset of news headlines, articles, and their categories for text classification. This disparity has hindered progress in machine learning and Natural Language Processing (NLP) research for these languages.
Most Indic languages share similarities, but they utilize different writing scripts, making it more challenging to accurately predict news categories. Since English is used widely around the world, and many researchers have studied how to classify text in English, it has resulted in a surplus of classification datasets. The same is not the case for Indic languages. Although there are existing datasets for Indian languages, they have some limitations, such as a lesser number of categories or inconsistent classification criteria for news articles, which narrows down the scope of research in this field. While the IndicNLP News Article dataset covers the major languages it is limited by the count of target labels and high accuracy Kakwani et al. (2020); Kulkarni et al. (2022). This calls for the creation of more intricate datasets to effectively assess the performance of models. In essence, there is a need for complex datasets that can thoroughly evaluate model effectiveness.
Text classification, a critical task in both machine learning and natural language processing (NLP), involves categorizing text documents into predefined classes based on their content. While there are a handful of publicly available datasets related to news, they often lack diversity in categories and sources, potentially leading to biased results.
Transformer-based models, like LongFormer [^1] https://huggingface.co/docs/transformers/model_doc/longformer Beltagy et al. (2020), require datasets with varying sequence lengths due to their sensitivity to text length. This highlights the need for specialized datasets to develop models for these low-resource languages. Hence, we introduce L3Cube-IndicNews, a comprehensive IndicNews Classification Dataset, sourced from diverse news websites specifically targeting low-resource languages.
This dataset encompasses over 3 lakh records, distributed across 12 diverse news categories, offering an extensive resource for supervised text classification. Each language dataset contains more than 26,000 rows, covering at least 10 significant news categories.
L3Cube-IndicNLP [^2] https://github.com/l3cube-pune/indic-nlp presents monolingual and multilingual models tailored to each Indian regional language. The repository includes individual BERT models for the languages focused on in this work. We conduct a comparative analysis of various monolingual and multilingual BERT models, including L3Cube monolingual BERT Joshi (2022a, b), monolingual SBERT Deode et al. (2023); Joshi et al. (2023a), IndicSBERT (multilingual) Deode et al. (2023) and IndicBERTKakwani et al. (2020).
The key contributions of this work are as follows:
  • Introduction of L3Cube-IndicNews, an extensive document classification dataset spanning ten significant Indian languages, each consistent with a range of 12 target labels. The dataset can also be used for news article headline-generation tasks.
  • The corpus comprises three sub-datasets (IndicNews- SHC, LPC, and LDC) catering to short, medium, and long documents, each with varying sentence lengths but consistent target labels.
  • The datasets are bench-marked using state-of-the-art pre-trained BERT models: L3Cube monolingual BERT, L3Cube monolingual SBERT, L3Cube IndicSBERT (multilingual) and IndicBERT.
2 Related work
IndicNLP News Article Classification333 https://github.com/AI4Bharat/indicnlp_corpus#indicnlp-news-article-classification-dataset dataset is part of the AI4Bharat-IndicNLP Dataset Kunchukuttan et al. (2020) that consists of news articles in 10 Indian languages categorized into classes like sports, entertainment, business, politics, and lifestyle. While this dataset contains a substantial number of records, it falls short in terms of the variety of categories available for news articles in each language, thereby limiting its diversity.
Multi Indic Languages News Dataset444 https://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-datasethttps://www.kaggle.com/datasets/shaz13/multi-indic-languages-news-dataset is a dataset publicly available on Kaggle. It is a multi-language news dataset from Times Internet for various Indian languages. This data contains columns named title, link, description, long_description, id. Despite the extensive size and diversity of this dataset, encompassing a rich collection of records across numerous Indian languages, it lacks language-wise and news category-wise segregation, which hampers clarity and ease of use.
Varta [^5] https://huggingface.co/datasets/rahular/varta is a multilingual dataset for headline generation. It encompasses 41.8 million news articles in 14 Indic languages and English. This data is sourced from DailyHunt. This dataset is well-organized and includes a large number of records for each language, covering all the minor details. However, this dataset contains only those articles written by DailyHunt’s partner publishers resulting in a biased nature towards a particular narrative or ideology that can affect the representativeness and diversity of the dataset. Aralikatte et al. (2023) From every language, they randomly sample 10,000 articles each for validation and testing. On average, Varta articles have 17 sentences, and the headlines have just over one sentence. A typical article sentence contains about 18 words, and a headline sentence contains 11 words. While the dataset is a large-scale, high-quality dataset for Indic languages, the headlines in this dataset are 39% smaller than the average sentence in an article.
iNLTK666 https://github.com/goru001/inltk is an openly accessible dataset primarily comprising data for 13 Indic languages, sourced from Wikipedia articles. It encompasses over 12,000 cleaned rows for each of these languages Arora (2020). This dataset comprises publicly accessible data for languages like Hindi, Bengali, Punjabi, Kannada, and Oriya. For languages such as Gujarati, Malayalam, Marathi, and Tamil, they have created their dataset by extracting information from Wikipedia articles. While the dataset boasts a substantial volume of records for each language, it falls short in terms of categorizing the data into specific news categories. Furthermore, in some languages, it includes not only news-related articles but also other types of content, leading to inefficiency and inconsistency in the dataset’s content and structure.
ACTSAMukku and Mamidi (2017) focuses on building a gold-standard annotated corpus of Telugu sentences to support Telugu Sentiment Analysis. The raw data is scraped from five different Telugu news websites viz. Andhrabhoomi, Andhrajyothi, Eenadu, Kridajyothi and Sakshi. In total, they collected over 453 news articles which were then filtered down to 321 articles relevant to their work.
Language Datasets available Categories present Articles
Hindi BBC Articles
India, International, Entertainment, Sports, Others
172K
BBC Hindi News Articles
India, Pakistan, News, International, Sntertainment, port, Science, China, Learningenglish, Social, Southasia, Business, Institutional, Multimedia
4335
Bengali AI4Bharat-IndicNLP Dataset
Entertainment, Sports
14K
Soham Articles
Kolkata, State, National, International, Sports, Entertainment
72K
Marathi AI4Bharat-IndicNLP Dataset
Entertainment, Lifestyle, Sports
4.5K
iNLTK Headlines
State, Entertainment, Sports
85K
Telugu AI4Bharat-IndicNLP Dataset
Entertainment, Business, Sports
24K
Tamil AI4Bharat-IndicNLP Dataset
Entertainment, Politics, Sports
11.7K
iNLTK Headlines
Tamil-cinema, Business, Spirituality
127K
Gujarati AI4Bharat-IndicNLP Dataset
Business, Entertainment, Sports
2K
iNLTK Headlines
Entertainment, Business, Tech
31K
Kannada AI4Bharat-IndicNLP Dataset
Entertainment, Lifestyle, Sports
30K
iNLTK-IndicNLP News Category
Entertainment, Sports, Tech
6.3K
Odia AI4Bharat-IndicNLP Dataset
Business, Crime, Entertainment, Sports
30K
iNLTK-IndicNLP News Category
Sports, Business, Entertainment
19K
Malayalam AI4Bharat-IndicNLP Dataset
Business, Entertainment, Sports, Technology
6K
iNLTK Headlines
Entertainment, Sports, Business
12K
Punjabi AI4Bharat-IndicNLP Dataset
Business, Crime, Entertainment, sports
3.1K
iNLTK -IndicNLP News Category
Politics, Non-politics
800
Table 1: Available Datasets in Indic languages
3 Curating the dataset
We introduce L3Cube-IndicNews, a comprehensive dataset compilation designed to facilitate the classification of both short text and long documents. Within IndicNews, there are three meticulously crafted supervised datasets: Short Headlines Classification (SHC), Long Document Classification (LDC), and Long Paragraph Classification (LPC).
Featuring a broad spectrum of information, the dataset comprises a minimum of 10 distinct categories corresponding to 10 prominent languages spoken in India. The careful organization of both language and newly defined categories enhances overall clarity. Sourced from multiple websites, it ensures a diverse array of content, exclusively from reputable news sources. Notably, there are no constraints on the length of articles and headlines, emphasizing a commitment to quality. In essence, it stands as a high-quality, versatile, and meticulously curated dataset.
3.1 Data collection
For the dataset curation process, we identified several websites to collect a substantial number of articles for each news category. The Hindi language dataset was scraped from Jansatta777 https://www.jansatta.com/. The Marathi language dataset was scraped from Lokmat888 https://www.lokmat.com/. The Bengali language dataset was scraped from Aajkal999 https://www.aajkaal.in/, Ganashakti101010 https://ganashakti.com/, BBC111111 https://www.bbc.com/bengali, Anandabazar121212 https://www.anandabazar.com/, abnews24131313 https://www.sangbadpratidin.in/, and Sangbadpratidin141414 https://www.abnews24.com/. The Telugu language dataset was scraped from ABP Telugu151515 https://telugu.abplive.com/. The Tamil Language dataset was scraped from Hindu Tamil161616 https://www.hindutamil.in/. The Gujarati language dataset was scraped from ABP Gujarati171717 https://gujarati.abplive.com/. The Kannada language dataset was scraped from Kannada Prabha181818 https://www.kannadaprabha.com/ and PublicTV191919 https://publictv.in/. The Odia language dataset was scraped from Odisha Bhaskar202020 https://odishabhaskar.com/ and Dharitri212121 https://www.dharitri.com/https://www.dharitri.com/ . The Malayalam language dataset was scraped from Madhyamam222222 https://www.madhyamam.com/. The Punjabi language dataset was scraped from Khabarwaale232323 https://www.khabarwaale.com/.
The data was gathered through the utilization of the urllib package for managing URL requests, coupled with the BeautifulSoup package for extracting data from the HTML content of the requested URL. Every website had organized its news articles into predefined categories, and during the scraping process, we retained this categorization to utilize it as the target label.
Each dataset was originally scraped to contain 3 columns: Title, Category, and News. Then 35% of the news article from column ‘News’ was used to create a fourth column ‘Sub article’, primarily containing a subset of news articles. The final curated dataset underwent shuffling and cleaning. We then divided these datasets into three supervised datasets: Short Headlines Classification (SHC), Long Document Classification (LDC), and Long Paragraph Classification (LPC).
Short Headlines Classification (SHC): This dataset contains the headlines of news articles paired with their respective categorical labels.
Long Paragraph Classification (LPC): In this dataset, each record contains a sub-article of news with its respective categorical label.
Long Document Classification (LDC): This dataset contains records having an entire news article with its corresponding categorical label.
3.2 Data Preprocessing
At first, we break down the article into sentences, making sure we handle the punctuation and sentence boundaries in Indic languages correctly. Then, we carefully break each sentence into tokens and filter them for special characters. We selectively retain tokens with an initial character aligned with the specific Indic language character set. We also created a dataset-cleansing function wherein each text is tokenized and then undesired elements, including words with specific characters, substrings, or minimal length are eliminated. Furthermore, regular expression was used to filter out words that don’t match the script of the language. This data refinement process ensures the dataset is free from unwanted elements, making it optimal for further use.
3.3 Dataset Statistics
Each dataset consists of more than 26,000 rows and 10-12 categories of news articles per dataset. Each of the 10 datasets was split into the train, test, and validation datasets in the ratio of 80:10:10.
Language Labels
Hindi (11)
Auto, Business, Crime, Education, Entertainment, Health, International, Nation, Politics, Sports, Technology
Bengali (10)
Sports, National, International, Kolkata, State, Politics, Entertainment, Technology, Editorial, Lifestyle
Marathi (12)
Auto, Bhakti, Crime, Education, Fashion, Health, International, Manoranjan, Politics, Sports, Tech, Travel
Telugu (10)
Business, Crime, Education, Entertainment, Jobs, Lifestyle, Politics, Sports, Technology, World
Tamil (10)
Auto, Business, Crime, Education, Entertainment, Health, India, Lifestyle, Politics, World
Gujarati (10)
Astro, Auto, Business, Crime, Entertainment, International, Nation, Sports, State, Technology
Kannada (10)
Business, Crime, Cuisine, Entertainment, International, Nation, Politics, Sports, State, Technology
Odia (10)
Lifestyle, Entertainment, News, Crime, Business, Health, Politics, Career, Sports, Editorial
Malayalam (10)
Crime, Entertainment, Gulf, International, Kerala, Lifestyle, National, Opinion, Sports, Technology
Punjabi (10)
World, Sports, Religious, Education, Transfer & Appointments, Literature, Health, National, Crime, Lifestyle
Table 2: Languages and their categorical labels
Language Train Test Validation Total
Hindi 30851 3835 3835 38521
Bengali 23970 2997 2996 29963
Marathi 22014 2761 2750 27525
Telugu 21103 2640 2650 26393
Tamil 25030 3129 3129 31288
Gujarati 26472 3347 3417 33236
Kannada 24642 3058 3058 30758
Odia 27420 3445 $34$ 34299
Malayalam 28000 3500 3500 35000
Punjabi 25494 3189 3186 31869
Table 3: Distribution of dataset into train, test, and validation in the ratio 80:10:10.
Figure 1: SHC Dataset Overview
Figure 2: LPC Dataset Overview
Figure 3: LDC Dataset Overview
4 Models
4.1 L3Cube Monolingual BERT242424 https://arxiv.org/abs/2211.11418 for Indic languages
We use the monolingual BERT models for the 10 Indic languages, released by L3cube-Pune252525 https://github.com/l3cube-pune as the base models. These models are termed as HindBERT262626 https://huggingface.co/l3cube-pune/hindi-bert-v2, BengaliBERT272727 https://huggingface.co/l3cube-pune/bengali-bert, MahaBERT282828 https://huggingface.co/l3cube-pune/marathi-bert, TeluguBERT292929 https://huggingface.co/l3cube-pune/telugu-bert, TamilBERT303030 https://huggingface.co/l3cube-pune/tamil-bert, GujaratiBERT313131 https://huggingface.co/l3cube-pune/gujarati-bert, KannadaBERT323232 https://huggingface.co/l3cube-pune/kannada-bert, OdiaBERT333333 https://huggingface.co/l3cube-pune/odia-bert, MalayalamBERT343434 https://huggingface.co/l3cube-pune/malayalam-bert, PunjabiBERT353535 https://huggingface.co/l3cube-pune/punjabi-bert. These models are fine-tuned on the existing multilingual models like MuRIL Khanuja et al. (2021), xlmRoBERTa Conneau et al. (2019), and IndicBERT on the monolingual corpus.
4.2 L3Cube Indic Sentence BERT [^36] https://arxiv.org/pdf/2304.11434.pdf models (Monolingual)
We also evaluate L3Cube monolingual Indic SBERT models that are HindSBERT373737 https://huggingface.co/l3cube-pune/hindi-sentence-bert-nliJoshi et al. (2023b), BengaliSBERT383838 https://huggingface.co/l3cube-pune/bengali-sentence-bert-nli, MahaSBERT [^39] https://huggingface.co/l3cube-pune/marathi-sentence-bert-nli, TeluguSBERT404040 https://huggingface.co/l3cube-pune/telugu-sentence-bert-nli, TamilSBERT414141 https://huggingface.co/l3cube-pune/tamil-sentence-bert-nli, GujaratiSBERT424242 https://huggingface.co/l3cube-pune/gujarati-sentence-bert-nli, KannadaSBERT434343 https://huggingface.co/l3cube-pune/kannada-sentence-bert-nli, OdiaSBERT444444 https://huggingface.co/l3cube-pune/odia-sentence-bert-nli, MalayalamSBERT454545 https://huggingface.co/l3cube-pune/malayalam-sentence-bert-nli, and PunjabiSBERT464646 https://huggingface.co/l3cube-pune/punjabi-sentence-bert-nli.
4.3 L3Cube Indic SBERT [^47] https://huggingface.co/l3cube-pune/indic-sentence-bert-nli (Multilingual)
IndicSBERT is the first multilingual SBERT model trained specifically for Indic languages. Sentence-BERT (SBERT) Reimers and Gurevych (2019) is a modified version of the BERT Devlin et al. (2018) architecture designed to generate sentence representations for the improved semantic similarity between sentences. The SBERT uses a Siamese network Koch et al. (2015) and is trained using specific datasets like STS, resulting in representations specifically geared for semantic similarity.
4.4 AI4Bharat indicBERT [^48] https://huggingface.co/ai4bharat/indic-bert
IndicBERT is a multi-lingual AlBERT model provided by AI4Bharat exclusively pre-trained in 12 Indian languages. It is pre-trained on AI4Bharat IndicNLP Corpora of around 9 billion tokens.
5 Evaluation
We evaluated each Indic dataset using the monolingual BERT model, multilingual BERT and SBERT models provided by L3Cube, and the indicBERT model provided by AI4Bharat. The results of evaluating these models on the curated datasets are shown in Table 4.
Language Model SHC LDC LPC
HindiBERT
86.600 91.681 88.097
Hindi
HindiSBERT
86.518
91.534
87.810
IndicSBERT
85.870
91.604
86.922
IndicBERT
83.910
86.809
84.424
BengaliBERT
82.549
95.293 85.195
Bengali
BengaliSBERT
82.749
94.141
84.470
IndicSBERT
80.674
92.457
85.232
IndicBERT
80.007
90.757
82.457
MarathiBERT
91.163 94.706
86.731
Marathi
MarathiSBERT
91.017
94.349
87.439
IndicSBERT
90.510
93.987
87.103
IndicBERT
89.388
92.627
85.222
TeluguBERT
89.810
92.765 91.818
Telugu
TeluguSBERT
90.416
92.651
91.098
IndicSBERT
87.916
92.348
91.515
IndicBERT
88.371
92.943
91.811
TamilBERT
81.785
84.521
81.300
Tamil
TamilSBERT
81.720
86.122
79.573
IndicSBERT
81.209
84.227
79.703
IndicBERT
81.275
83.226
80.571
GujaratiBERT
89.898 95.278 90.640
Gujarati
GujaratiSBERT
89.808
95.060
90.091
IndicSBERT
88.613
95.277
89.573
IndicBERT
87.909
90.854
89.050
KannadaBERT
91.410
94.706
87.704
Kannada
KannadaSBERT
89.340
94.539
87.345
IndicSBERT
90.255
94.833
87.835
IndicBERT
87.965
91.857
87.358
OdiaBERT
83.399
92.940
84.557
Odia
OdiaSBERT
84.137 93.424
83.918
IndicSBERT
82.065
93.054
84.441
IndicBERT
82.661
86.962
82.322
MalayalamBERT
80.440 88.573 81.705
Malayalam
MalayalamSBERT
80.171
88.201
81.171
IndicSBERT
77.780
88.029
79.672
IndicBERT
75.114
83.098
76.457
PunjabiBERT
85.456
90.182
84.163
Punjabi
PunjabiSBERT
90.363
94.866
90.333
IndicSBERT
89.490
94.224
88.446
IndicBERT
91.725
94.781
90.358
Table 4: Accuracies for all the models trained on SHC, LDC, and LPC datasets in percentage (%)
Language Model SHC LDC LPC
Hindi
HindiSBERT
86.209
88.451
88.279
Bengali
BengaliSBERT
82.615
93.126
89.289
Marathi
MarathiSBERT
91.017
94.349
87.439
Telugu
TeluguSBERT
88.674
92.310
91.174
Tamil
TamilSBERT
81.754
84.902
80.993
Gujarati
GujaratiSBERT
88.792
92.378
90.609
Kannada
KannadaSBERT
89.797
94.212
88.260
Odia
OdiaSBERT
84.137
93.424
83.918
Malayalam
MalayalamSBERT
79.114
87.743
80.428
Punjabi
PunjabiSBERT
94.744
93.614
93.782
Table 5: Accuracies for monolingual SBERT model trained on mixed dataset (SHC + LPC + LDC) in percentage (%)
6 Results
For evaluating the models on these datasets, we use accuracy as our main evaluation metric to understand the performance of our models. Table 4 presents the accuracies obtained by fine-tuning our models on the datasets. The confusion matrices for the Kannada dataset trained and tested on SHC, LDC, and LPC respectively are illustrated in Figures 4, 5, and 6.
Key observations are outlined as follows:
  • In most cases, the L3Cube Monolingual model tends to exhibit superior performance in terms of accuracy, as demonstrated by the metrics presented in the table across all corpora.
  • Among the Short Headlines Classification (SHC), Long Document Classification (LDC), and Long Paragraph Classification (LPC) datasets, LDC demonstrated the most impressive results after fine-tuning text classification. This aligns with expectations, given that long document data inherently contains more information compared to shorter-length datasets.
  • On the other hand, SHC reported comparatively lower accuracy scores across all three document types. This may be attributed to the fact that news headlines can sometimes encompass more generalized information, potentially leading to some degree of confusion for the models.
Figure 4: Confusion matrix for the Kannada SHC dataset
Figure 5: Confusion matrix for the Kannada LDC dataset
Figure 6: Confusion matrix for the Kannada LPC dataset
Table 5 presents the accuracies obtained by training the L3Cube Monolingual Sentence BERT model on a mix of all the 3 datasets (SHC + LPC + LDC) and then evaluating them on individual datasets. A cross-analysis was conducted by evaluating the performance of L3Cube Monolingual Indic Sentence BERT model on individual test datasets. Upon fine-tuning the Monolingual SBERT model on a mixed dataset, notable results were observed, with the Long Document Classification (LDC) task exhibiting the most impressive performance.
7 Future Work
There is potential for expanding the dataset’s labels to achieve broader category coverage or facilitate further category expansion. The dataset creation process has been streamlined through automation, eliminating the need for manual typing. Looking forward, there is a prospect of implementing the curation of a manually typed or verified dataset, ensuring an even higher level of accuracy and reliability for future applications.
Currently, we have achieved proficiency in Long Document Classification (LDC) for a single language, and the performance in Long Paragraph Classification (LPC) is satisfactory. However, we acknowledge that the performance in Short Headlines Classification (SHC) falls short of optimization. To address this, we have implemented a unified model selection, specifically opting for the monolingual SBERT based on its significant performance on individual datasets for each language.
This chosen model exhibits competence in handling mixed datasets, comprising of SHC, LPC, and LDC, and its performance has been thoroughly examined. It’s important to note that our current focus remains on a single language. Looking ahead, we recognize the potential for future developments, particularly in the realm of cross-dataset analysis.
8 Conclusion
In this research paper, we introduce L3Cube-IndicNews, a comprehensive collection of three labeled datasets, encompassing over 3 lakh records in ten different Indic languages. These datasets are designed for text classification tasks in the context of Indian languages. Within this paper, we provide an in-depth overview of the creation process, which involves the use of 10 to 12 distinct categorical labels to curate these supervised datasets. To assess the effectiveness of these datasets, we conducted fine-tuning on BERT-based models, serving as a valuable benchmark for future research and development.
Our experiments involved four key models: L3Cube Monolingual BERT, L3Cube Monolingual SBERT, L3Cube-IndicSBERT (multilingual), and AI4BHARAT IndicBERT. Notably, our findings indicate that the BengaliBERT Model achieved the highest accuracy when applied to the LDC dataset. We believe that the availability of our datasets will contribute significantly to the enhancement of NLP support for the Indic languages, promoting its growth and development in this field.
9 Acknowledgements
This work was carried out under the mentorship of L3Cube, Pune. We would like to express our gratitude towards our mentor, for his continuous support and encouragement. This work is a part of the L3Cube-IndicNLP Project.